# Learning From Software Failures: A Case Study at a National Space Research Center

### v2

@@ Abstract.
To address this gap, we conducted a case study through 10 in-depth interviews with research software engineers at a national space research center. We examine how they learn from failures: how they gather, document, share, and apply lessons from failures. To assess the transferability of our findings, we include data from 5 additional interviews at other HROs. Our findings provide insight on how software engineers learn from failures in practice. To summarize our findings: (1) failure learning is informal, ad-hoc, and inconsistently integrated into SDLC; (2) recurring failures persist due to the absence of structured processes; and (3) key challenges, including time constraints, knowledge loss due to team turnover & fragmented documentation, and weak process enforcement, undermine efforts to systematically learn from failures. Our findings contribute to a deeper understanding of how software engineers learn from failures and offer guidance for improving failure management practices.
copyright: noneconference: 48th International Conference on Software Engineering; April 12–18, 2026; Rio de Janeiro, Brazil
@@ 1. Introduction
A core engineering principle is the systematic analysis of failures to prevent their recurrence (Petroski, 1994; Dalcher, 1994). This approach has been effective across various engineering disciplines, contributing to the low failure rates of e.g., buildings, aircraft, and nuclear reactors. A key method for institutionalizing learning from software failures is the use of postmortem and retrospective practices, where engineers systematically analyze failures and gather lessons learned (Collier et al., 1996). However, despite the recognized importance of these practices, they are rarely followed consistently as part of the software engineering process within organizations (Vieira et al., 2019; Kasi et al., 2008; Dingsøyr, 2005; Glass, 2002; Keegan and Turner, 2001). We believe that one contributing factor to the persistent high rate of software failures is the limited adoption and effectiveness of practices to learn from failures (Anandayuvaraj and Davis, 2023). Although existing works have provided general methods for conducting postmortems (Dingsøyr, 2005; Tiedeman, 1990; Collier et al., 1996; Myllyaho et al., 2004; Stålhane et al., 2003; Dingsøyr et al., 2001) and observations on how postmortem practices have been implemented (Desouza et al., 2005; Washburn Jr et al., 2016; Birk et al., 2002; Kasi et al., 2008; Lyytinen and Robey, 1999; Dingsøyr et al., 2007; Vieira et al., 2019), we lack empirical insight on the practices and challenges faced by engineers to learn from failures: how they gather, document, share, and apply lessons from failures..
Such insights might be particularly valuable to study in high-reliability organizations (HROs), such as those in aerospace, healthcare, and finance, because their software systems operate under strict safety, security, and performance constraints to maintain reliability (Christianson et al., 2011). In these environments, software failures can have severe consequences, as demonstrated by recent software incidents with catastrophic impacts (Anandayuvaraj et al., 2024; Ko et al., 2014; Wong et al., 2017, 2009; Zollers et al., 2005; National Research Council, 2007), ranging from financial losses to threats to human life. As highlighted by the National Academy of Engineers (National Research Council, 2007), software engineering failures are widespread, with annual costs estimated in the trillions of dollars (Krigsman, 2009), emphasizing the need to reduce software failures. Consequently, organizations should invest resources into software failure analysis and learning processes to reduce recurrence and improve reliability.
@@
Our results provide insight into how software engineers in HROs learn from failures. Although HROs are expected to excel at failure management, our study reveals that structured processes for learning from failures is often informal and ad hoc. Systemic barriers, such as time constraints, documentation gaps, and team turnover, limit effective learning. As a result, recurring failures persist despite general process adherence. In summary, we make the following contributions:
@@
    We identify key challenges — technical, organizational, and cultural — that impact learning from failures.
@@ 2.1. Learning from failures in engineering
An essential principle of engineering is to analyze failures, learn from them, and implement mitigation to reduce the risk of recurrence in the future (Petroski, 1994; Reason, 1990; Dalcher, 1994). This principle has been successful in many engineering disciplines, where lessons from failures has contributed to the low failure rates of, e.g., medical devices, aircraft, railway systems, and nuclear reactors. For example, there are databases of failures in medical devices (Commissioner, 2022), aviation (ASR, [n. d.]), aerospace (NASA, 2023), railways (Rai, 2017), and chemicals (Inv, [n. d.]), to enable practitioners in those fields to learn from past failures. Methodologies like Six Sigma are used by manufacturing (Gijo et al., 2011), chemical (Falcón et al., 2012), civil (Morales et al., 2016), and automotive (Antony et al., 2005) engineering fields to systematically analyze errors and improve processes (Antony, 2002).
@@
Figure 1. The general process of learning from engineering failures (Dingsøyr, 2005; Tiedeman, 1990; Collier et al., 1996; Myllyaho et al., 2004; Stålhane et al., 2003; Dingsøyr et al., 2001; Kasi et al., 2008). The dotted region indicates our study focus, annotated with the associated RQs.
@@
As a starting point, a few organizations have shared practices and experience reports. Practices are often structured around (1) gathering, (2) documenting, (3) sharing, and (4) applying lessons from failures (Kasi et al., 2008), as illustrated in Figure 1. Structured reviews or retrospectives are recommended to (1) gather lessons from technical and socio-organizational factors that contributed to failures (Dingsøyr, 2005; Beyer et al., 2016; NAS, [n. d.]a). For example, Google describes a blameless postmortem approach to foster a culture of learning (Beyer et al., 2016). Documentation practices are recommended to (2) document not only root causes but also broader context, decision rationale, mitigation efforts, and lessons in an accessible and durable format (Dingsøyr, 2005; Beyer et al., 2016; NAS, [n. d.]b, [n. d.]a; Johnson, 2000). For example, organizations like NASA (NAS, [n. d.]b) and ESA (Les, 2024) maintain databases of lessons learned. (3) Sharing often relies on both formal mechanisms like on-boarding training, cross-team reviews, and newsletters, as well as informal mechanisms like mentoring and reading clubs (Dingsøyr, 2005; Beyer et al., 2016). However, literature on (4) applying lessons learned is limited and their long-term effectiveness remains uncertain (Pope and Sillito, 2024; Sillito and Pope, 2024b). A suggested practice is to create action items from lessons (Dingsøyr, 2005; Beyer et al., 2016).
However, many other organizations lack processes to study and learn from failures (Vieira et al., 2019; Kasi et al., 2008; Dingsøyr, 2005; Glass, 2002; Keegan and Turner, 2001). In addition, there is skepticism about how carefully these processes are followed in organizations that have them (Kasi et al., 2008; Sillito and Pope, 2024a). The aforementioned exemplar organizations publish their policies but rarely share engineers’ experiences in following them, nor whether it leads to sustained learning, especially beyond the postmortem room into the rest of the engineering process. To address this gap, we turn to HROs, which are theoretically expected to exemplify continuous learning and failure management (Christianson et al., 2011).
3. Reseach Questions
In our analysis of the literature discussed in section 2, we observed that there is little empirical data on how software engineers learn from software failures: how they gather, document, share, and apply lessons from failures. We lack in-depth qualitative descriptions of industry experiences, creating a gap in our understanding of practices to learn from failures. We do not know which how far practice is from theory, which practices are effective, nor the challenges in this engineering process. To fill this gap, we ask:
@@ 4. Methodology
Figure 2. Study methodology. We situated our work as a case study at a space research center, complemented by pilot data from other HROs. Several sources informed our interview protocol. Two author-analysts analyzed the data.
@@ 4.2. Case study setting: A space-related HRO
As a national space research center responsible for supporting mission-critical projects, it exemplifies characteristics of an HRO (Christianson et al., 2011). This department operates in a high-risk context and shows sensitivity to operational failures. Its structure and mission mandate a commitment to resilience and deferring to technical expertise. The department’s role in software development for space missions, along with its responsibility for software quality assurance, aligns with the expectations of an HRO that must prioritize reliability, continuous learning, and failure management. Given these properties, we hypothesized that the organization would exhibit practices for learning from software failures, making it an appropriate site to conduct our study.
@@ 4.3. Protocol design and development
To develop the initial interview protocol, we studied the background and related literature (section 2), which provided a solid foundation of existing knowledge on learning from failures and highlighted gaps that informed our research questions and subsequently our interview questions. We identified knowledge gaps and created the initial protocol, structured along typical postmortem and failure management process recommendations.
We refined the protocol through practice interviews (2 subjects) and pilot interviews (5 subjects). For practice, the lead author interviewed two graduate students with prior SE work experience to iteratively improve the interview questions. These practice interviews led to modifications across our interview protocol to improve depth and focus. In the pilot, we conducted 5 interviews with practitioners from various HRO industries. The subjects for the pilot study were recruited using convenience sampling, drawing from two groups: part-time professional master’s students at our university (pilot interviews 1–3) and practitioners from the lead author’s internship at another organization (pilot interviews 4–5).
During these pilot studies we gathered insight about how these practitioners and their organizations learn from failures. We include data from the pilot study where feasible to evaluate transferability of our findings from the case study at the center to other organizations. The pilot study interviews led to modifications to improve focus on failure recurrence, communication, and automation. In the Reflective Example section, new follow-up questions were added to explore whether issues reoccurred after initial fixes and what actions were taken to prevent recurrence. The Postmortem Analysis section was expanded to examine when issues are shared across teams and stakeholders. Additionally, a new Automation section was introduced to investigate desired tools for documenting and learning from failures. These changes aimed to capture deeper insights into long-term learning and prevention strategies.
@@ 4.4. Data collection
Recruitment: As a summer intern, the lead author facilitated interviews within the organization. The lead author’s internship advisors, two of the secondary authors, provided a list of potential subjects for the study. These two authors were also case study interview subjects. [^1] To eliminate bias, those authors-cum-subjects did not participate in data analysis. 9 other candidates were invited to participate in the study, out of which 8 accepted. All candidates were provided with a consent form which included brief questions on demographics (5 questions) and development practices (4 questions).
@@
Interviews: All interviews were conducted by the lead author, via Zoom. The median interview duration was 53 minutes.
@@ 4.5. Data analysis
Figure 3. Saturation curve. Interviews are plotted in the order in which they were conducted. Pilot interviews were conducted before the case study; they are separated by a dash-ed black line. Each interview covered many topics in detail (red line, blue line). However, as the green line shows, our results saturated—i.e., we stopped observing new perspectives—around interview 8.
@@ 4.6. Threats to validity
Internal Threats concern factors that may influence the validity of cause-and-effect interpretations. As with any qualitative research, data interpretation carries an inherent risk of subjectivity. To enhance the reliability of our findings, we employed an iterative approach to develop the codebook, involving a collaborative approach with multiple raters reviewing, discussing, and developing the codes, and achieved a high level of agreement. We recognize the potential for bias since the case study was conducted at an organization where one of the authors was employed as an intern.
@@ 5. Results and Analysis
We organize our results into three themes that answer our research questions. Our analysis reveals both 6 new findings and 5 findings that have been partially discussed in prior works. The challenges faced by this organization to learn from failures mirror barriers in conducting postmortems identified in prior works: perceived low benefit, lack of time, funding, culture, procedure, and mechanisms for failure learning, and the tacit and situated nature of failure knowledge (Kasi et al., 2008). We also observed additional challenges, including fragmented knowledge, difficulty recalling lessons, overconfidence in general practices, shifting responsibility to individuals (especially newcomers), concerns about bureaucratic processes, reliance on informal SDLCs, and knowledge loss from turnover. A summary of results is outlined in Table 3.
Table 3. Findings from our thematic analysis mapped to research questions. The Lit. column notes prior works discussing related findings.
RQ Section and Summary Lit.
RQ1 (section 5.1.1) Lesson gathering is ad hoc, informal, and individually driven; reflection is not routine.  (Kasi et al., 2008; Vieira et al., 2019)
(section 5.1.2) Documentation practices are inconsistent, leaving failure knowledge fragmented & largely tacit.  (Kasi et al., 2008; Vieira et al., 2019)
RQ2 (section 5.1.3) Lessons are shared via informal conversations & senior mentorship within teams. Cross-project sharing is rare.  (Kasi et al., 2008; Lyytinen and Robey, 1999)
(section 5.1.3) Applying past lessons is driven by individual memory or initiative. Improvements are not part of a structured cycle.
(section 5.2.3) Some examples of ad hoc learning exist, but they are exceptions and not integrated into broader practice.  (Sillito and Pope, 2024a)
RQ3 (section 5.2.1) Recurring failures are common due to lack of structured processes for reflection and improvement.
(section 5.2.2) Teams often assume good practices suffice to prevent failures; this masks systemic gaps and limits formal learning.
(section 5.3.1) Time and resource constraints limit opportunities for failure learning; structured approaches are seen as burdensome.  (Kasi et al., 2008; Desouza et al., 2005; Vieira et al., 2019; Lyytinen and Robey, 1999; Dingsøyr, 2005)
@@
(section 5.3.3) Knowledge is often lost due to team turnover and poor documentation of lessons learned.
(section 5.3.4) Failure knowledge is fragmented across tools; weak enforcement and siloed teams hinder accessibility and application.
5.1. Theme 1: Failure learning is ad hoc

### v3

@@ Abstract.
To address this gap, we conducted a case study through 10 in-depth interviews with research software engineers at a national space research center. We examine how they learn from failures: how they gather, document, share, and apply lessons learned. To assess the transferability of our findings, we include data from 5 additional interviews at other HROs. Our findings provide insight on how software engineers learn from failures in practice. To summarize our findings: (1) failure learning is informal, ad-hoc, and inconsistently integrated into SDLC; (2) recurring failures persist due to the absence of structured processes; and (3) key challenges, including time constraints, knowledge loss due to team turnover & fragmented documentation, and weak process enforcement, undermine efforts to systematically learn from failures. Our findings contribute to a deeper understanding of how software engineers learn from failures and offer guidance for improving failure management practices.
Postmortem, Retrospective, Failure Analysis, Software Engineering
journalyear: 2026copyright: rightsretainedconference: 2026 IEEE/ACM 48th International Conference on Software Engineering; April 12–18, 2026; Rio de Janeiro, Brazilbooktitle: 2026 IEEE/ACM 48th International Conference on Software Engineering (ICSE ’26), April 12–18, 2026, Rio de Janeiro, Brazildoi: 10.1145/3744916.3773149isbn: 979-8-4007-2025-3/26/04
@@ 1. Introduction
Engineers expect some defects (Kuutila et al., 2020; Costello, 1984) but strive to eliminate severe ones that cause incidents: undesired, unplanned, software-induced events causing substantial loss (Leveson, 1995; Avizienis et al., 2004). The presence of these defects, whether caught internally or as incidents, signifies a failure in the software engineering process. Understanding these failures is vital for failure mode analysis (Reifer, 1979; Song, 2012; Ishimatsu et al., 2010) to help identify and prioritize risks (Fairbanks, 2010).
A core engineering principle is the systematic analysis of failures to prevent their recurrence (Petroski, 1994; Dalcher, 1994). This approach has been effective across various engineering disciplines, contributing to the low failure rates of e.g., buildings, aircraft, and nuclear reactors. Postmortem and retrospective practices, where engineers systematically analyze failures and gather lessons learned (Collier et al., 1996), are among the recommended methods for institutionalizing learning from software failures. However, despite the recognized importance of these practices, they are rarely followed consistently as part of the software engineering process within organizations (Vieira et al., 2019; Kasi et al., 2008; Dingsøyr, 2005; Glass, 2002; Keegan and Turner, 2001). We believe that one contributing factor to the persistent high rate of software failures is the limited adoption and effectiveness of practices to learn from failures (Anandayuvaraj and Davis, 2023). Although existing works have provided general methods for conducting postmortems (Dingsøyr, 2005; Tiedeman, 1990; Collier et al., 1996; Myllyaho et al., 2004; Stålhane et al., 2003; Dingsøyr et al., 2001) and observations on how postmortem practices have been implemented (Desouza et al., 2005; Washburn Jr et al., 2016; Birk et al., 2002; Kasi et al., 2008; Lyytinen and Robey, 1999; Dingsøyr et al., 2007; Vieira et al., 2019), we lack empirical insight on the practices and challenges faced by engineers to learn from failures: how they gather, document, share, and apply lessons learned..
Such insights are particularly valuable to study in high-reliability organizations (HROs), such as those in aerospace, healthcare, and finance, because their software systems operate under strict safety, security, and performance constraints to maintain reliability (Christianson et al., 2011). In these environments, software failures can have severe consequences, as demonstrated by recent software incidents with catastrophic impacts (Anandayuvaraj et al., 2024; Ko et al., 2014; Wong et al., 2017, 2009; Zollers et al., 2005; National Research Council, 2007), ranging from financial losses to threats to human life. As highlighted by the National Academy of Engineers (National Research Council, 2007), software engineering failures are widespread, with annual costs estimated in the trillions of dollars (Krigsman, 2009), emphasizing the need to reduce software failures. Consequently, organizations should invest resources into software failure analysis and learning processes to reduce recurrence and improve reliability.
@@
Our results provide insight into how software engineers in HROs learn from failures. Although HROs are expected to excel at failure management, our study reveals that structured processes for learning from failures is often informal and ad hoc. Systemic barriers, such as time constraints, documentation gaps, and team turnover, limit effective learning. As a result, recurring failures persist despite general process adherence. In summary, our contributions are:
@@
    We identify key challenges – technical, organizational, and cultural – that impact learning from failures.
@@ 2.1. Learning from Failures in Engineering
An essential principle of engineering is to analyze failures, learn from them, and implement mitigation to reduce risk of future recurrence  (Petroski, 1994; Reason, 1990; Dalcher, 1994). This principle has been successful in many engineering disciplines, where lessons from failures has contributed to the low failure rates of, e.g., medical devices, aircraft, railways, and nuclear reactors. For example, there are databases of failures in medical devices (Commissioner, 2022), aviation (ASR, [n. d.]), aerospace (NASA, 2023), railways (Rai, 2017), and chemicals (Inv, [n. d.]), to enable practitioners in those fields to learn from past failures. Methodologies like Six Sigma are used by manufacturing (Gijo et al., 2011), chemical (Falcón et al., 2012), civil (Morales et al., 2016), and automotive (Antony et al., 2005) engineering fields to systematically analyze errors and improve processes (Antony, 2002).
@@
Figure 1. The general process of learning from engineering failures (e.g., vulnerabilities, incidents) (Dingsøyr, 2005; Tiedeman, 1990; Collier et al., 1996; Myllyaho et al., 2004; Stålhane et al., 2003; Dingsøyr et al., 2001; Kasi et al., 2008). The dotted region indicates our study focus, annotated with the associated RQs.
@@
As a starting point, a few organizations have shared practices and experience reports on learning from failures. Practices are often structured around (1) gathering, (2) documenting, (3) sharing, and (4) applying lessons from failures (Kasi et al., 2008), as illustrated in Figure 1. Structured reviews or retrospectives are recommended to (1) gather lessons from technical and socio-organizational factors that contributed to failures (Dingsøyr, 2005; Beyer et al., 2016; NAS, [n. d.]a). For example, Google describes a blameless postmortem approach to foster a culture of learning (Beyer et al., 2016). Documentation practices are recommended to (2) document not only root causes but also broader context, decision rationale, mitigation efforts, and lessons in an accessible and durable format (Dingsøyr, 2005; Beyer et al., 2016; NAS, [n. d.]b, [n. d.]a; Johnson, 2000). For example, organizations like NASA (NAS, [n. d.]b) and ESA (Les, 2024) maintain databases of lessons learned. (3) Sharing often relies on both formal mechanisms like on-boarding training, cross-team reviews, and newsletters, as well as informal mechanisms like mentoring and reading clubs (Dingsøyr, 2005; Beyer et al., 2016). However, literature on (4) applying lessons learned is limited and their long-term effectiveness remains uncertain (Pope and Sillito, 2024; Sillito and Pope, 2024b; Anandayuvaraj et al., 2023). A suggested practice is to create action items from lessons (Dingsøyr, 2005; Beyer et al., 2016).
However, many organizations lack processes to study and learn from failures (Vieira et al., 2019; Kasi et al., 2008; Dingsøyr, 2005; Glass, 2002; Keegan and Turner, 2001). In addition, there is skepticism about how well these processes are followed in organizations that have them (Kasi et al., 2008; Sillito and Pope, 2024a). The aforementioned exemplar organizations publish their policies, but rarely share engineers’ experiences with them, nor whether it leads to sustained learning, especially beyond the postmortem room into the rest of the engineering process. To address this gap, we turn to HROs, which are theoretically expected to exemplify continuous learning and failure management (Christianson et al., 2011).
3. Research Questions
In our analysis of the literature discussed in section 2, we observed that there is little empirical data on how software engineers learn from software failures: how they gather, document, share, and apply lessons from failures. We lack in-depth qualitative descriptions of industry experiences, creating a gap in our understanding of practices to learn from failures. We do not know how far practice is from theory, which practices are effective, nor the challenges in this engineering process. To fill this gap, we ask:
@@ 4. Methodology
Figure 2. Study methodology. We situated our work as a case study at a space research center, complemented by pilot data from other HROs. Two author-analysts analyzed the data.
@@ 4.2. Case Study Setting: A Space-related HRO
As a national space research center responsible for supporting mission-critical projects, it exemplifies characteristics of an HRO (Christianson et al., 2011). The department’s role in software development for space missions, along with its responsibility for software quality assurance, aligns with the expectations of an HRO that must prioritize reliability, continuous learning, and failure management. Given these properties, we hypothesized that the organization would exhibit practices for learning from software failures, making it an appropriate site to conduct our study.
The engineers in the department occupy a hybrid role: they are responsible for developing and maintaining mission-critical software, while also engaging in research activities such as experimenting with new programming languages and fault-tolerance techniques. In a broad sense, they can be defined as Research Software Engineers (RSEs) (Goth et al., 2025; Sims, 2021; Segal, 2005), who research and develop production-grade software, but are distinct from RSEs who primarily develop scientific analysis software.
@@ 4.3. Protocol Design and Development
To develop the initial interview protocol, we studied the related literature (section 2), which provided a foundation of existing knowledge on learning from failures and highlighted gaps that informed our research questions and subsequently our interview questions. We identified knowledge gaps and created the initial protocol, structured along typical postmortem and failure management processes.
We refined the protocol through practice interviews (2 subjects) and pilot interviews (5 subjects). For practice, the lead author interviewed two graduate students with prior SE work experience to iteratively improve the interview questions. These practice interviews led to modifications across our interview protocol to improve depth and focus. In the pilot, we conducted 5 interviews with practitioners from various HRO industries. During these pilot studies we gathered insight about how these practitioners and their organizations learn from failures. We include data from the pilot study where feasible to evaluate transferability of our findings from the case study at the center to other organizations. The pilot study interviews led to modifications to improve focus on failure recurrence, communication, and automation. In the Reflective Example section, new follow-up questions were added to explore whether issues reoccurred after initial fixes and what actions were taken to prevent recurrence. The Postmortem Analysis section was expanded to examine when issues are shared across teams and stakeholders. Additionally, a new Automation section was introduced to investigate desired tools for documenting and learning from failures. These changes aimed to capture deeper insights into long-term learning and prevention strategies.
@@ 4.4. Data Collection
Recruitment: For the case study, the lead author facilitated interviews within the organization as a summer intern. The lead author’s internship advisors, two of the secondary authors, provided a list of potential subjects for the study. These two authors were also interviewed for the study. [^1] To eliminate bias, these authors-cum-subjects did not participate in data collection, coding, or analysis, except to provide feedback as member checking. A total of 10 interviews were conducted within the case study: 2 with these co-author subjects and 8 with other engineers. In addition, as stated in section 4.3, we asses transferability of findings from the case study with the 5 pilot interviews. The subjects for the pilot study were recruited using convenience sampling, drawing from two groups: part-time professional master’s students at our university (pilot interviews 1–3) and practitioners from the lead author’s internship at another organization (pilot interviews 4–5). All candidates were provided with a consent form which included brief questions on demographics (5 questions) and development practices (4 questions).
@@
Interviews: All interviews were conducted in English by the lead author, via Zoom. The median interview duration was 53 minutes.
@@ 4.5. Data Analysis
Figure 3. Saturation curve. Interviews are plotted in the order in which they were conducted. Pilot interviews were conducted before the case study; they are separated by a dash-ed black line. Each interview covered many topics in detail (red line, blue line). However, as the green line shows, our results saturated–i.e., we stopped observing new perspectives–around interview 8.
@@ 4.6. Threats to Validity
Internal Threats concern factors that may influence the validity of cause-and-effect interpretations. As with any qualitative research, data interpretation carries an inherent risk of subjectivity. To enhance the reliability of our findings, we iteratively developed the codebook, involving a collaborative approach with multiple raters reviewing, discussing, and developing the codes, and achieved a high level of agreement. We recognize the potential for bias since the case study was conducted at an organization where the authors were employed. For member checking (Birt et al., 2016), two senior subjects (who were co-authors but not involved in the analysis) reviewed our findings to validate our interpretations.
@@ 5. Results and Analysis
We organize our results into three themes that answer our research questions. A summary of results is outlined in Table 3. Our analysis reveals 6 new findings and 5 findings that have been partially discussed in prior works.
Table 3. Findings from our thematic analysis mapped to research questions. The Lit. column notes prior works discussing related findings. The Pilot Sub. column indicates pilot subjects whose interviews support transferability of the finding.
RQ Section and Summary Lit. Pilot Sub.
RQ1 (section 5.1.1) Lesson gathering is ad hoc, informal, & individually driven; reflection is not routine.  (Kasi et al., 2008; Vieira et al., 2019) 1, 2, 3, T
(section 5.1.2) Documentation practices are inconsistent, leaving failure knowledge fragmented & largely tacit.  (Kasi et al., 2008; Vieira et al., 2019) 1, 2, 3, T
RQ2 (section 5.1.3) Lessons are shared via informal conversations & senior mentorship within teams. Cross-project sharing is rare.  (Kasi et al., 2008; Lyytinen and Robey, 1999) 1, 2, 3, 4
(section 5.1.3) Applying past lessons is driven by individual memory or initiative. Improvements aren’t part of structured cycle. 1, 2, 4, T
(section 5.2.3) Some examples of ad hoc learning exist, but they are exceptions & not integrated into broader practice.  (Sillito and Pope, 2024a) 1, 2, 4, T
RQ3 (section 5.2.1) Recurring failures are common due to lack of structured processes for reflection & improvement. 1, 3, 4, T
(section 5.2.2) Teams often assume good practices suffice to prevent failures: masks systemic gaps & limits formal learning. 1, 2, 4
(section 5.3.1) Time & resource constraints limit opportunities for failure learning; structured methods are seen as burdensome. (Kasi et al., 2008; Desouza et al., 2005; Vieira et al., 2019; Lyytinen and Robey, 1999; Dingsøyr, 2005) 1, 2, 3, 4, T
@@
(section 5.3.3) Knowledge is often lost due to team turnover & poor documentation of lessons learned. 1, T
(section 5.3.4) Failure knowledge is fragmented across platforms, hindering accessibility & application. 1, 2, 4, T
5.1. Theme 1: Failure Learning is Ad–hoc